# Zero-Shot Statistical Tests for LLM-Generated Text Detection using Finite Sample Concentration Inequalities

### v2

@@ 4.3 Statistical test for detection between an LLM and human
5 Preliminary experiments
@@
We conduct our numerical analysis on the following pre-trained language models: GPT-2 small, GPT-2 medium, GPT-2 large, GPT-2 XL, and GPT-Neo. Through experiments with different generative and evaluator models, we examine whether the log-perplexity of a short portion of text converges to the average cross-entropy. Our experiments measure these values across generated text and analyze their performance over different configurations. Our setup includes generating tokens with pre-trained models and recording each token’s selection probability and calculated metrics.
5.1 Same generative and evaluator model
@@
In the first set of experiments, we employ GPT-2 to generate a series of $100$ tokens, beginning with the fixed prompt "Jack". We use the model’s conditional probability distribution for each token generation step to sample the next token. Note that for the white-box model of GPT-2, probability distributions are accessible. We calculate each generated token’s empirical entropy and log-perplexity and repeat this process for comparisons. We use Softmax-normalized probabilities to select the next token and store the generated token and its probability distribution. For each sub-string of length $N$ starting from the first token in the generated sequence, we compute the log-perplexity $l_{A}(\mathbf{Y}_{N})$, and the empirical entropy $h_{N}(A,A)(\mathbf{Y}_{N})$. The results are shown in Figures (1(a)-1(d)). We consistently observe that the numerical results confirm Theorem 1 that the log-perplexity converges to the average entropy when the generative and evaluator models are the same.
@@
5.2 Different generative and evaluator models
To extend our numerical analysis to the case with different generative and evaluator models, we generate a string using the following generative models: GPT-2 medium, GPT-2 large, and GPT-2 XL. Then, we calculate the log-perplexity of these strings using the evaluator model GPT2-small. We calculate the cross-entropy of the strings under each generative model and the evaluator model (GPT2- small). The results are shown in Figures (2(a)-2(c)). Results in these figures confirm Theorem 2. In particular, we observe that when the evaluator and generative models are different, the log-perplexity of the string converges to the average cross-entropy of the string under generative and evaluator models.
@@ A.8 Proof of Proposition 3
Appendix B Related literature
Preemptive detection methods. One detection approach is to record (Krishna et al., 2024) or watermark (Kirchenbauer et al., 2023) all generated text. The method proposed in Kirchenbauer et al. (2023) takes the last generated token in the prefix and uses it to seed an RNG, which randomly places $50\%$ of the possible subsequent tokens in a green list and the remaining in a red list. During sampling, the algorithm boosts the probability of sampling a green word, resulting in a higher than $50\%$ of the final text consisting of green words. For a survey on watermarking methods, we refer readers to Amrit and Singh (2022). It is noteworthy to mention that preemptive approaches can only be applied if the LLM owner cooperates by allowing full control over the generative model.
Unlike preemptive detection methods, post-hoc detection methods can be used without needing cooperation from model developers. The two main categories in this group are neural-based and statistical-based methods.
@@ Appendix B Related literature
Neural-based methods. Significant efforts have been made to develop neural-based methods (e.g., features-based classifiers). However, a key concern with this approach is poor robustness, e.g., against ambiguous semantics (Schaaff et al. (2023)) and the neural-based method’s limited ability to detect LLM-generated misinformation (Schuster et al. (2020)). Recent efforts have been enhancing training methodology (e.g., Kumarage et al., 2023; Tu et al., 2023). However, these methods provide only heuristics without theoretical analysis to guarantee that results would hold irrespective of the specific features of their studied texts. Additionally, there are limitations for the feature-based classifiers, including challenges in training models with the hype in developing new LLMs, an increase in the variety of topics and writing styles, and legal concerns associated with training on human data, such as privacy concerns.
Statistical-based methods. Most research on statistics-based methods focuses on white-box statistics, black-box statistics, and linguistics feature statistics. The majority of works in this category also only provide heuristics.
  • White-Box Statistics. The existing white-box detection methods primarily apply logit-based statistics or perturbed-based methods. Closer to our work are logit-based statistics. In the logit-based stream, Log-likelihood is one of the most widely used measures (Solaiman et al. (2019)). Other measures include using GTLR based on rank-likelihood (Gehrmann et al. (2019)), the Log-likelihood Ratio Ranking (LRR) proposed by Su et al. (2023), entropy and Kullback-Leibler (KL) divergence (e.g., Lavergne et al. (2008)), and perplexity (e.g., Verma et al., 2021; Vasilatos et al., 2023). The majority of research in applying white-box statistics proposes heuristics.
  • Black-Box Statistics. Based on the observation that human-generated texts require more revisions when LLMs are tasked with editing the text compared to LLM-generated texts, several works, including Yang et al., 2023; Zhu et al., 2023; Mao et al., 2024 employ LLMs to continue writing truncated texts under review and differentiate between human and LLM texts by calculating the n-gram similarity between the continuation and the original text. Quidwai et al. (2023) proposes calculating cosine similarity to distinguish between the two. While beneficial in that they do not require access to the logits of the source model, black-box statistics methods are challenging due to the substantial overhead of accessing the LLM and long response times.
  • Linguistics Features Statistics. Several recent studies on detecting LLM-generated text have proposed detection methodologies based on linguistics feature statistics. For a survey on features, we refer readers to Beresneva (2016); for a summary of recent research using linguistic features for AI detection, we refer readers to Wu et al. (2023). One limitation of applying linguistic feature statistics is that their availability relies heavily on extensive corpus statistics and various types of LLMs. Our solution of deriving a theoretical bound that guarantees achieving the required level of detection certainty when a small corpus is available can resolve this hurdle.
@@
Possibility of detection & Robustness. Robust detection methods are being developed, and increasingly sophisticated evasion methods are being devised to circumvent these detectors, creating an ongoing contest between detection and evasion. The evasion of detection methods can be through (i) prompt engineering by general users who change the writing style of LLM-generated text or (ii) through paraphrasing attacks by adversaries. Recent research examines the ability of different detection methods against the mentioned evasion methods, including, e.g., Sadasivan et al. 2023; Krishna et al. 2024; Zhang et al. 2024. As a theoretical work in this stream, Chakraborty et al. (2023) provides (detection) possibility results for detecting machine-generated and human-generated texts. Specifically, results characterize the number of samples for the likelihood-ratio-based detector to achieve an AUROC of $\epsilon$. In binary classification, the ROC Curve is a graphical representation that illustrates the performance of a binary classification model at various thresholds, and the AUROC (Area Under the ROC Curve) quantifies model’s ability to distinguish between classes. As the decision threshold changes, AUROC shows the trade-off between the True Positive Rate (TPR) and the False Positive Rate (FPR). For a survey on possibility/impossibility of detecting AI-generated text, please refer to Ghosal et al. (2023).

### v3

@@ 4.3 Statistical test for detection between an LLM and human
5 Experiments
In this section, we present two main sets of experiments: (i) experiments with white-box access to the conditional probability distributions for token generation to examine our theoretical results, and (ii) experiments that relax the white-box access assumption. In the black-box setting, we also examine robustness to decoding strategies and adversarial attacks.
Figure 2: Overview of the experiments
Lastly, we provide experiments to demonstrate that the log-perplexity under the evaluator model and the average cross-entropy between the generative and the evaluator model converge.
5.1 White-box detection
Datasets. We use the datasets used in Mitchell et al. (2023) for our experiments. That includes the news articles from the XSum dataset (Narayan et al. 2018), prompted stories from Reddit’s WritingPrompts dataset (Fan et al. 2018), and Wikipedia paragraphs from the SQuAD contexts (Rajpurkar 2016) that contain human-written questions and machine-generated answers from passages from Wikipedia articles. Each experiment uses 150 to 500 examples for evaluation. We prompt the first 30 tokens of the actual text to the LLM to generate machine-generated text and obtain the generated text. For QA datasets, we prompt the question. In all experiments, a positive signal is that the focal LLM generates text, and all experiments use an equal number of positive and negative examples. The language models we examine for whitebox detection are LLAMA 3 (8B parameters), GPT-NEOX Erebus (20B parameters), and QWEN (32B parameters). We use AUROC and TPR @ FPR=1%, which follows the recent research that uses this evaluation paradigm (Krishna et al. 2024; Hans et al. 2024; Soto et al. 2024) as our performance metrics.
AUROC. The Area Under the Receiver Operating Characteristic (ROC) Curve is a standard widely adopted detection performance measure, e.g., Verma et al. 2023; Mitchell et al. 2023, and we use this metric consistent with the literature. The ROC curve plots the True Positive Rate (TPR) against the False Positive Rate (FPR) at all possible classification thresholds, showing the tradeoff between correctly detecting LLM-generated text and mistakenly flagging human-written text.
TPR @ FPR=1%. Results of experiments in the literature, e.g., in Dugan et al. 2024, suggest that the false positive rates of detectors can be dangerously high, which makes them unfit for regulatory actions. To show the validity of our theoretical results, we examine the accuracy captured by the TPR while maintaining the low FPR.
Benchmarks. We compare the performance of our method with various existing zero-shot detection methods that similarly leverage the predicted token-wise conditional distributions of the source model. The benchmarks we use apply detection tests based on token log-probabilities, token (log-)ranks, likelihood curvature, or (a measure of) predictive entropy (Gehrmann et al. 2019; Solaiman et al. 2019; Ippolito et al. 2019; Mitchell et al. 2023; Hans et al. 2024). Detection based on token log-probabilities uses the source model’s average token-wise log-probability. Detection based on token (log-)rank uses the average observed (log-)rank of the tokens in the passage based on the model’s conditional distributions. The entropy-based detection benchmark inspired by Gehrmann et al. 2019 and applied by Mitchell et al. 2023 hypothesizes that machine-generated texts are more "in-distribution" for the model, leading to lower entropy predictive distributions. Detection based on Mitchell et al. 2023’s framework states that a machine-generated text is sampled mostly at the mode of its text distribution, while human texts might lay anywhere on the text distribution. Lastly, Hans et al. (2024) applies a normalized measure of perplexity by dividing log-perplexity by average cross-entropy.
5.2 White-box detection results
AUROC. In these experiments, model samples are generated by sampling from the raw conditional distribution with temperature 1.
XSum SQuAD WritingPrompts
Method LLAMA 3 8B GPT-NEOX Erebus QWEN 32B LLAMA 3 8B GPT-NEOX Erebus QWEN 32B LLAMA 3 8B GPT-NEOX Erebus QWEN 32B
$\log p(x)$ 0.99* 0.84 0.99* 0.91 0.75 0.60 1.00 0.95 1.00*
Rank 0.75 0.69 0.69 0.71 0.71 0.55 0.82 0.80 0.81
LogRank 0.99* 0.87 0.99* 0.93* 0.81 0.62 1.00 0.97* 1.00*
Entropy 0.39 0.70 0.40 0.37 0.66 0.70* 0.04 0.36 0.02
DetectGPT 0.78 0.95 0.99 * 0.55 0.78 0.62 0.68 0.97* 0.99
Binoculars 0.78 0.95 0.99 * 0.55 0.78 0.62 0.68 0.97* 0.99
Ours 0.99 0.99 1.00 0.99 0.99 0.97 0.98* 1.00 1.00
Diff (zero-shot) 0.00 0.04 0.01 0.06 0.18 0.27 -0.02 0.03 0.00
Roberta (base) 0.98* 0.95 0.92* 0.97* 0.92 0.69* 0.97* 0.95* 0.74*
Roberta (large) 0.98* 0.98* 0.92* 0.95 0.93* 0.68 0.96 0.93 0.65
Ours 0.99 0.99 1.00 0.99 0.99 0.97 0.98 1.00 1.00
Diff (supervised) 0.01 0.01 0.08 0.02 0.06 0.28 0.01 0.05 0.26
Table 1: AUROC for detecting samples from the given models on the datasets using our and other baseline methods. The best AUROC values are in bold, and the second-best values are marked with an asterisk (*). The rows Diff(zero-shot) and Diff(supervised) show our AUROC improvement over the strongest zero-shot and supervised baseline methods.
Our method achieves its highest improvement in detecting LLM-generated text for SQuAD, increasing AUROC from 0.04 to 0.27. It also improves detection for XSum (0.015 AUROC) and WritingPrompts (0.025 AUROC), though with a smaller margin. An explanation for the weaker performance of existing detectors on SQuAD is its higher entropy variance (0.128) compared to XSum (0.045) and WritingPrompts (0.074). Since our method relies on the difference between log-perplexity and empirical entropy, it performs more effectively when entropy variance is high while aligning with likelihood-based methods when entropy is less variable. Finally, our method significantly outperforms existing likelihood-based approaches in detecting Qwen-generated text for SQuAD.
TPR @ FPR=1%. In Table 2, we show the accuracy captured by TPR, while the FPR is maintained low (1%).
XSum SQuAD WritingPrompts
Method LLAMA 3 8B GPT-NEOX Erebus QWEN 32B LLAMA 3 8B GPT-NEOX Erebus QWEN 32B LLAMA 3 8B GPT-NEOX Erebus QWEN 32B
$\log p(x)$ 0.89 0.02 0.99* 0.01 0.05 0.01 1.00 0.46 1.00
Rank 0.22 0.06 0.20 0.07 0.18 0.01 0.49 0.39 0.43
LogRank 0.99 0.05 0.99* 0.02 0.06 0.01 1.00 0.54 1.00
Entropy 0.02 0.09 0.01 0.01 0.04 0.00 0.00 0.00 0.00
Binoculars 0.99 0.89* 0.99* 0.96 0.86* 0.26* 0.97* 0.99 0.99
Ours 0.97* 1.00 1.00 0.96 0.92 0.72 0.90 0.99 1.00
Diff (zero-shot) -0.02 0.11 0.01 0.00 0.06 0.46 -0.10 0.00 0.00
Table 2: TPR at FPR=0.01. The best TPR values are in bold, and the second-best values are marked with an asterisk (*). The row Diff(zero-shot) shows our TPR improvement over the strongest zero-shot baseline methods.
The experiment’s results show the notable superiority of our method in fact-based datasets (e.g., SQuAD), where the human-written and LLM-generated texts tend to be similar, resulting in lower performance of conventional detection methods. Specifically, we observe the maximum gap between ours and the second-best methods in the SQuAD dataset with QWEN 32B as the evaluator. Even when the FPR is held low, the human and LLM texts are more distinguishable in datasets with less fact-based and more creative content. Also, smaller language models (e.g., LLAMA 2 8B) are easier to detect.
5.3 Black-box detection
Datasets. We use the datasets in Dugan et al. 2024 that consist of data in 8 domains of abstracts, books, news, poetry, recipes, Reddit, reviews, and Wikipedia. The non-adversarial dataset consists of 14,971 human-written from publicly-available pre-2022 datasets and 509,014 LLM-generated documents for a total of 6,287,820 texts when including adversarial attacks.
Benchmarks. We compare the performance of our method with the detectors in Dugan et al. 2024 that include neural detectors, zero-shot detectors, and commercial detectors. The neural detectors are RoBERTa-B (GPT-2), RoBERTa-L (GPT-2), RoBERTa-B (ChatGPT), and RADAR. The zero-shot detectors are GLTR, Fast DetectGPT, Binoculars, and LLMDet. The commercial detectors are GPTZero, Originality, Winston, and ZeroGPT. In this section, we do not have white-box access to any of the detectors, and we use tiiuae/falcon-rw-1B as our scoring model.
Repetition penalty. This penalty down-weighs the probability of tokens that previously appeared in the context window by some multiplicative factor $\theta$, resulting in less repetitive output. Following Keskar et al. 2019, we use $\theta=1.2$ for our experiments.
Adversarial attacks. When selecting adversarial attacks, consistent with Dugan et al. 2024, we assume that the adversary uses predefined transformations or heuristics that degrade typical detection cues, different from the gradient-based methods. Accordingly, we include the following 8 black-box, query-free attacks:
  1. 1.
    Alternative Spelling (AS): uses British spelling instead of the American spelling.
  2. 2.
    Article Deletion (AD): deletes "the", "a", "an."
  3. 3.
[…]
  4. 4.
    Number Swap (NS): randomly shuffles number digits.
  5. 5.
    Misspelling (MS): inserts common misspellings.
  6. 6.
    Synonym Swap (SS): swaps tokens with highly similar BERT (Devlin et al. (2019)) candidate tokens.
  7. 7.
    Upper Lower Swap (ULS): swaps the case of words.
  8. 8.
    Whitespace Addition (WSA): adds spaces between characters.
Note that two other adversarial attacks, i.e., Homoglyph and Zero-width space, are possible. Homoglyph swaps characters in a text with visually similar yet technically different characters, often from other alphabets or Unicode blocks. Zero-width space inserts a Unicode character (often U+200B) that does not occupy visible space. Dugan et al. (2024)’s dataset for these attacks was incomplete, not allowing us to provide results for these attacks. However, these two attacks are easy to remove with the following pre-processing:
  • Removing or normalizing invisible characters like zero-width spaces
  • Mapping suspicious non-ASCII characters to standard ASCII or a canonical form
5.4 Black-box detection results
Accuracy at FPR=5%. Consistent with Dugan et al. (2024), we evaluate accuracy by TPR while fixing FPR at 5%. In the experiments, the open-source chat models include Llama-c, Mistral-c, and MPT-c. The open-source non-chat models include Mistral, MPT, and GPT2. The closed-sourced chat models include c-GPT, GPT4, and Cohere, and the closed-sourced non-chat models include Cohere and GPT3. For the repetition penalty, we set the multiplicative factor as $\theta=1.2$.
Open-Source Closed-Source
Chat Models Non-Chat Models Chat Models Non-Chat Models
Rep. Penalty? × × × ×
GPTZero 98.4 82.5 9.4 4.8 88.5 53.4
Originality 97.7 72.5 89.0 51.2 89.0 85.4
Winston 96.6 78.3 29.5 11.3 93.7 68.1
ZeroGPT(*) 90.5 54.9 16.0 0.3 65.8 72.7
R-B GPT2 77.9 26.2 60.5 35.4 41.7 52.5
R-L GPT2 71.4 19.5 67.2 53.4 34.7 48.6
R-B CGPT 75.0 39.3 14.9 1.7 38.1 39.0
RADAR 85.6 66.4 48.3 31.8 75.3 67.7
GLTR 83.9 38.3 44.5 0.5 54.3 63.7
F-DetectGPT 96.2 40.5 79.7 0.6 74.1 86.3
LLMDet 47.5 16.5 38.4 3.7 18.5 32.9
Binoculars 99.7 60.6 72.4 0.6 92.1 95.0
Ours 97.6 60.7 51.7 71.2 68.8 86.5
Table 3: Accuracy Score at FPR=5% for all detectors across model groups and sampling strategies. Asterisks (*) indicate that the detector was unable to achieve the target FPR. For accuracy of 67%-100%, we highlight green, 33%-66% we highlight yellow, and for 0%-32% we highlight red.
For open-source chat models without a repetition penalty, our method achieves an accuracy of 97.6%, indicating that when text maintains its natural stylistic and structural cues, our detector can accurately flag LLM-generated content. For open-source non-chat models, the addition of a repetition penalty improves performance, with accuracy rising from 51.7% without the penalty to 71.2% with the penalty. In the closed-source setting, our detector again performs robustly—68.8% for chat models and a high 86.5% for non-chat models, demonstrating our method’s capacity to generalize across different text styles. Overall, these findings indicate that our detector is particularly strong when the underlying text generation preserves the natural language features that our method is designed to leverage, and that applying a repetition penalty can further enhance accuracy in contexts where repetitive patterns might otherwise obscure these signals.
Adversarial attacks
None AS AD IP NS MS SYN ULS WSA
RoB-B GPT2 59.1 55.6 37.1 56.9 55.9 43.8 71.5 18.8 45.2
RoB-L GPT2 56.7 52.4 33.2 55.1 51.7 39.5 79.4 19.3 40.1
RoB-B CGPT 44.8 43.3 38.0 5.2 44.3 42.1 39.6 31.7 0.1
RADAR 70.9 70.8 67.9 73.7 71.0 69.5 67.5 70.4 66.1
GPTZero 66.5 64.9 61.0 66.2 65.8 65.1 61.0 56.5 66.2
ZeroGPT 65.5 65.4 59.7 64.9 64.7 64.7 18.8 54.5 64.2
Originality 85.0 83.6 71.4 85.1 86.0 78.6 96.5 75.8 84.9
Winston 71.0 68.9 66.9 69.8 69.0 67.5 63.6 56.8 46.8
GLTR 62.6 61.2 52.1 61.4 59.9 59.8 31.2 48.1 45.8
F-DGPT 73.6 71.6 64.7 72.0 68.2 70.7 34.0 60.4 64.4
LLMDet 35.0 33.9 27.4 27.2 33.8 32.7 27.3 23.4 4.4
Binoculars 79.6 78.2 74.3 71.7 77.1 78.0 43.5 73.8 70.1
Ours 84.3 79.9 73.1 80.7 73.3 78.1 48.0 72.9 57.1
Table 4: Accuracy Score at FPR=5% for all detectors across different adversarial attacks. Abbrevations are: AS: Alternative Spelling, AD: Article Deletion, IP: Insert Paragraphs, NS: Number Swap, MS: Misspelling, SYN: Synonym Swap, ULS: Upper Lower Swap, WSA: Whitespace Addition. Cell colors: Red < 33, Yellow 33–66, Green $\geq$ 67.
Our detector maintains relatively high accuracy across multiple adversarial scenarios. In the baseline “None” condition, it records an accuracy of 84.3%. Under attacks such as Alternative Spelling (79.9%), Article Deletion (73.1%), Insert Paragraphs (80.7%), Number Swap (73.3%), Misspelling (78.1%), and Upper–Lower Swap (72.9%), the accuracy remains in the 70–80% range. However, similar to other zero-shot detectors, our performance is reduced under Synonym Swap (48.0%) and Whitespace Addition (57.1%). This detection performance drop is predictable because zero-shot detection methods often rely on fixed token distributions. Synonym substitutions replace surface-level tokens with semantically equivalent but lexically distinct items, thereby shifting the token distribution away from the detector models’ reference distribution. Whitespace manipulations alter standard sub-word segmentation by inserting or removing token boundaries, creating a discrepancy between the text’s actual token frequency and what the detector expects.
5.5 Convergence results
@@
We conduct our numerical analysis for convergence on the following pre-trained language models: GPT-2 small, GPT-2 medium, GPT-2 large, GPT-2 XL, and GPT-Neo. Through experiments with different generative and evaluator models, we examine whether the log-perplexity of a short portion of text converges to the average cross-entropy. Our experiments measure these values across generated text and analyze their performance over different configurations. Our setup includes generating tokens with pre-trained models and recording each token’s selection probability and calculated metrics.
@@
Same generative and evaluator model. In the first set of experiments, we employ GPT-2 to generate a series of $100$ tokens, beginning with the fixed prompt "Jack". We use the model’s conditional probability distribution for each token generation step to sample the next token. Note that for the white-box model of GPT-2, probability distributions are accessible. We calculate each generated token’s empirical entropy and log-perplexity and repeat this process for comparisons. We use Softmax-normalized probabilities to select the next token and store the generated token and its probability distribution. For each sub-string of length $N$ starting from the first token in the generated sequence, we compute the log-perplexity $l_{A}(\mathbf{Y}_{N})$, and the empirical entropy $h_{N}(A,A)(\mathbf{Y}_{N})$. The results are shown in Figures (3(a)-3(d)). We consistently observe that the numerical results confirm Theorem 1 that the log-perplexity converges to the average entropy when the generative and evaluator models are the same.
@@
Different generative and evaluator models. To extend our numerical analysis to the case with different generative and evaluator models, we generate a string using the following generative models: GPT-2 medium, GPT-2 large, and GPT-2 XL. Then, we calculate the log-perplexity of these strings using the evaluator model GPT2-small. We calculate the cross-entropy of the strings under each generative model and the evaluator model (GPT2- small). The results are shown in Figures (4(a)-4(c)). Results in these figures confirm Theorem 2. In particular, we observe that when the evaluator and generative models are different, the log-perplexity of the string converges to the average cross-entropy of the string under generative and evaluator models.
@@ A.8 Proof of Proposition 3
Appendix B Related work
@@
Supervised detection methods. Supervised detection methods learn to differentiate between human-written and LLM-generated text through labeled examples. For example, GPT2 Detector (Solaiman et al. 2019) fine-tunes RoBERTa on the output of GPT2, and ChatGPT Detector (Guo et al. 2023) fine-tunes RoBERTa on the HC3 dataset in Guo et al. 2023. Many other supervised detection methods also exist, including those that leverage neural representations (e.g., Bakhtin et al. 2019; Uchendu et al. 2020; Fagni et al. 2021) and those that leverage bag-of-words (Fagni et al. 2021). Although some supervised detection methods show high performance, it has been observed that they tend to overfit their training domains and source models (Bakhtin et al. 2019; Uchendu et al. 2020). A key concern with neural-based methods (e.g., feature-based classifiers) is their poor robustness, e.g., against ambiguous semantics (Schaaff et al. (2023)) and their limited ability to detect LLM-generated misinformation (Schuster et al. (2020)). Recent efforts have been enhancing training methodology (e.g., Kumarage et al., 2023; Tu et al., 2023). However, these methods provide heuristics without theoretical analysis to guarantee that results would hold irrespective of the specific features of their studied texts. Additionally, there are limitations for the feature-based classifiers, including challenges in training models with the hype in developing new LLMs, an increase in the variety of topics and writing styles, and legal concerns associated with training on human data, such as privacy concerns.
Zero-shot detection methods. Another stream of work attempts to distinguish machine-generated from human-written texts using statistics-based methods. Most research on statistics-based methods focuses on white-box statistics, black-box statistics, and linguistics feature statistics. Among those, white-box detection methods are closely related to our work. The existing white-box detection methods primarily apply logit-based statistics or perturbed-based methods. Closer to our work are logit-based statistics. In the logit-based stream, Log-likelihood is one of the most widely used measures (Solaiman et al. 2019). Other measures include using GTLR based on rank-likelihood (Gehrmann et al. 2019), the Log-likelihood Ratio Ranking (LRR) proposed by Su et al. (2023), entropy and Kullback-Leibler (KL) divergence (Lavergne et al. 2008), and perplexity ( Vasilatos et al. 2023; Wang et al. 2023b). Mitchell et al. (2023) observes that machine-generated texts tend to lie in the local curvature of the log probability and proposes DetectGPT, which has demonstrated reliable performance but has some limitations. Du et al. analyzes the limitations of DetectGPT, highlights DetectGPT’s need for computing many perturbations, making it a computationally intensive algorithm, and proposes targeted masking strategies (rather than random masking) to improve DetectGPT’s performance. Hans et al. (2024) proposes Binoculars that use a ratio of perplexity measurement and cross-perplexity, a notion of how surprising the next token predictions of one model are to another model. Like supervised detection methods, most statistics-based detection methods provide only heuristics.
Watermarking. One detection approach is to record (Krishna et al., 2024) or watermark (Kirchenbauer et al., 2023) all generated text. The method proposed in Kirchenbauer et al. (2023) takes the last generated token in the prefix and uses it to seed an RNG, which randomly places $50\%$ of the possible subsequent tokens in a green list and the remaining in a red list. During sampling, the algorithm boosts the probability of sampling a green word, resulting in a higher than $50\%$ of the final text consisting of green words. Note that watermarking requires the cooperation of the generating party/LLM owner to implement the green listing-red listing algorithm. For a survey on watermarking methods, we refer readers to Amrit and Singh (2022).
@@
Possibility of detection and robustness to attacks. Robust detection methods are being developed, and increasingly sophisticated evasion methods are being devised to circumvent these detectors, creating an ongoing contest between detection and evasion. The evasion of detection methods can be through (i) prompt engineering by general users who change the writing style of LLM-generated text or (ii) through paraphrasing attacks by adversaries. Recent research examines the ability of different detection methods against the mentioned evasion methods, including, e.g., Sadasivan et al. 2023; Krishna et al. 2024; Zhang et al. 2024. As a theoretical work in this stream, Chakraborty et al. (2023) provides (detection) possibility results for detecting machine-generated and human-generated texts. Specifically, results characterize the number of samples for the likelihood-ratio-based detector to achieve an AUROC of $\epsilon$. In binary classification, the ROC Curve is a graphical representation that illustrates the performance of a binary classification model at various thresholds, and the AUROC (Area Under the ROC Curve) quantifies model’s ability to distinguish between classes. As the decision threshold changes, AUROC shows the trade-off between the True Positive Rate (TPR) and the False Positive Rate (FPR). For a survey on possibility/impossibility of detecting AI-generated text, please refer to Ghosal et al. (2023).
Authorship attribution. The literature on human-written text author attribution is extensive and, similar to the recent literature on LLM text detection and attribution, has seen developments over time, e.g., use of statistical hypothesis testing versus discriminative methods like support vector machines, neural networks, etc. In author attribution with only human authors, there is no access to the text generation "model," and the attribution is only based on the samples of texts written by different human authors. For surveys on the different methods for human author attribution, we refer readers to Juola et al. 2008; Stamatatos 2009; Koppel et al. 2009, and for a comparison of different methods, we refer readers to Grieve (2007).
With the advent of LLMs, human author attribution extends to (1) human-written vs. machine-generated detection, (2) machine-generated vs. another machine-generated attribution, and (3) attribution in text generated through LLM and human collaboration. These shifts necessitate transitioning from human-centric stylometry to detection frameworks tailored for LLMs. For a survey that connects the older literature on authorship attribution to modern problems in the era of LLMs, we refer readers to Huang et al. (2025).
Statistical tests on LLM-generated text. Recent efforts have provided statistical tests for detecting LLM-generated text and determining whether cloud implementation works the same as the reference model. For example, Li et al. leverages the watermark key sequence and next token probabilities (NPTs) to introduce statistical tests for detecting LLM-generated text when the detection approach is watermarking, and their method applies to short and long text. Gao et al. (2024) introduces a framework for detecting whether an API provider has modified a language model’s output distribution without informing users. The paper formalizes the problem as a two-sample statistical testing task, where users collect samples from an API and compare them to a reference model’s outputs. They propose a hypothesis testing method using Maximum Mean Discrepancy (MMD) with a string kernel, first proposed in Gretton et al. (2012), to detect whether the API serves the expected model or a modified version.